Use the explicit formula an=a1+(n-1)d to find the 500th term of the sequence below.
24,31,38,45,52,..

Answer by nerdybill(7384)

Use the explicit formula an=a1+(n-1)d to find the 500th term of the sequence below.
24,31,38,45,52,..
an is the value of the nth term
a1 is the first term 24
n is the term number of interest 500
d is the difference 7
.
an = 24+(500-1)(7)
an = 24+(499)(7)
an = 24+3493
an = 3517 (value of the 500th term)